# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 2. Jordan type functions for $\mathbb{G}$-modules
$(\mathcal{U}_{\mathbb{G},r})_{*}:k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\quad\to\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$
the map of group algebras over $k[V_{r}(\mathbb{G})]$ induced by $\mathcal{U}_{\mathbb{G},r}$. In other words, $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the $k[V_{r}(\mathbb{G})]$ dual of the map on coordinate algebras $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$.
We readily observe that a closed embedding $i:\mathbb{H}\hookrightarrow\mathbb{G}$ of affine group schemes of finite type over $k$ determines the commutative square
(2.8.2) ${\mathbb{G}_{a(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{H},r})_{*}}$${\mathbb{H}_{(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}].}$
We recall and reformulate the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}$ for $\mathbb{G}_{(r)}$-modules of [14, Defn 2.1].
@@
where $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the map of “group algebras” for the map $\mathcal{U}_{\mathbb{G},r}$ of group schemes defined in Definition 2.8.
In [14, Ex 2.6], the reader will find a few explicit computations of $\Theta_{\mathbb{G},r}$. We mention two such examples consistent with the indexing used in this paper.
@@
Consider $\mathbb{G}=\mathbb{G}_{a}^{\times r}$ whose Lie algebra is the dimension $r$ Lie algebra $\mathfrak{g}_{a}^{\oplus r}$ (which is commutative, with $p$-th power map trivial). Identify $V_{1}(\mathbb{G}_{a}^{\times r})$ with $\mathbb{A}^{r}=\operatorname{Spec}\nolimits k[t_{0},\ldots,t_{r-1}]$ and $k(\mathbb{G}_{a}^{\times r})_{!}$ with $k[x_{0},\ldots,x_{r-1}]/(x_{i}^{p})$. Then
$\Theta_{\mathbb{G}_{a}^{\times r},1}\ =\ \sum_{i=0}^{r-1}t_{i}\otimes x_{i}\ \in\ k[V_{1}(\mathbb{G}_{a}^{\times r}]\otimes\mathfrak{u}(\mathfrak{g}_{a}^{\oplus r}).$
Consider $\mathbb{G}=\mathbb{G}_{a(r)}$ whose Lie algebra is $\mathfrak{g}_{a}$. Identify $V_{r}(\mathbb{G}_{(r)})$ with $\operatorname{Spec}\nolimits k[T_{0},\ldots,T_{r-1}]$ with $T^{i}$ given grading $p^{i}$ and $k\mathbb{G}_{(r)}$ with $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. Then
$\Theta_{\mathbb{G}_{a},r}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ +\ \cdots\quad\in\ k[V_{r}(\mathbb{G}_{a})]\otimes k\mathbb{G}_{a(r)},$
where the non-explicit, remaining terms of $\Theta_{\mathbb{G}_{a},r}$ each consist of a polynomial homogeneous of weighted degree $p^{r-1}$ in the $T_{i}$’s tensor a product of two or more $u_{i}$’s.
@@
    Let $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ be a closed subgroup scheme determining $\phi^{*}:k[V_{r}(\mathbb{G})]\twoheadrightarrow k[V_{r}(\mathbb{H})]$ and $\phi_{(r)*}:k\mathbb{H}_{(r)}\hookrightarrow k\mathbb{G}_{(r)}$. Then the restriction of $(\phi^{*}\otimes 1)(\Theta_{\mathbb{G},r})\in k[V_{r}(\mathbb{H})]\otimes k\mathbb{G}_{(r)}$ equals the image of $(1\otimes\phi_{(r)*})(\Theta_{\mathbb{H},r})$.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathbb{G},r}$ determines a $p$-nilpotent, $k[V_{r}(\mathbb{G})]$-linear operator $\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\ \to\ k[V_{r}(\mathbb{G})]\otimes M$.
@@
Consequently, the restriction of $\Theta_{\mathbb{G},r}=\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})$.
The naturality of $\mathbb{G}\mapsto\Theta_{\mathbb{G},r}$ with respect to closed embeddings $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ as in assertion (2) follows from the commutativity of (2.8.2) as verified in [14, Prop 2.9].
@@
The fact that this action is $p$-nilpotent follows immediately from the observation that the $i$-th power of $\Theta_{\mathbb{G},r}$ equals $(\mathcal{U}_{\mathbb{G},r})_{*}((1\otimes u_{r-1})^{i})\ =\ ((\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1}))^{i}$. Assertion (4) follows directly from assertion (1).
The last assertion is stated and proved in [14, Prop 2.11]. The proof proceeds by first observing that it suffices to verify that the restriction of the $k[V_{r}(\mathbb{G})]$-linear map $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ to $1\otimes k[\mathbb{G}_{(r)}]$ reads off the coefficient of $t^{p^{r-1}}$ in $k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\simeq k[V_{r}(\mathbb{G})]\otimes k[t]/t^{p^{r}}$. The authors then verify that the actions of $\mathbb{A}^{1}$ on $\mathbb{G}_{a(r)}$ and $V_{r}(\mathbb{G})$ are appropriately compatible. ∎
The following is an immediate corollary of Theorem 2.11.
Corollary 2.12.
The Jordan type of the $k(x)\otimes k[u]/u^{p}$-module obtained as the restriction of the $k(x)\otimes k\mathbb{G}_{(r)}$-module $k(x)\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}$ (as in Definiton 2.3) equals the Jordan type of  $k(x)\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r,M}\ \in\ End_{k(x)}(k(x)\otimes M)$.
For any graded $R$-module $M_{*}$, we denote by $(M_{*})^{\sim}$ the associated quasi-coherent sheaf on $\operatorname{Proj}\nolimits(R)$. We shall often consider graded $R$-modules of the form $R\otimes V$ where $V$ is a $k$-vector space and $R\otimes V$ inherits the grading of $R$.
If $M_{*}$ is a graded $R$-module, we denote by $M_{*}(n)$ the graded module obtained from $M_{*}$ by “shifting down by $n$” the degrees of $M_{*}$; thus, $(M_{*}(n))_{d}=M_{n+d}$. We use $(M_{*})^{\sim}(n)$ to denote $(M_{*}(n))^{\sim}$. The operator $\Theta_{\mathbb{G},r,M}$ of Theorem 2.11(3) multiplies degrees by $p^{r-1}$, so that viewed as an operator on graded modules one should write $\Theta_{\mathbb{G},r,M}$ as
@@
Because $k$ has positive characteristic, the informative survey by M. Beltrametti and L. Robbiano in [2] does not apply to our consideration of $\operatorname{Proj}\nolimits V_{r}(\mathbb{G})$. Because $V_{r}(\mathbb{G})$ is usually not generated by homogeneous elements of degree 1, we supplement the clear exposition of R. Hartshorne in [16, §II.5] with [8, Ex 9.5] when considering the relationship between graded modules over a graded, commutative $k$-algebra $R$ and sheaves of modules on $\operatorname{Proj}\nolimits(R)$. We also point out that our weighted projective spaces are not obtained by taking a quotient by an action of $\mathbb{G}_{m}$.
@@
Consider a polynomial algebra $k[T_{0},\ldots,T_{N}]$ with each $T_{i}$ given a “weighted” degree $d_{i}$. Assume that $I_{Y}\subset k[T_{0},\ldots,T_{N}]$ is an ideal generated by polynomials $F_{j}(\underline{T})$ which are (weighted) homogenous. Then, the zero locus $Y=Z(I_{Y})\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$ is a weighted projective variety associated to the graded algebra $k[T_{0},\ldots,T_{N}]/I_{Y}$.
Let $d$ be a positive integer divisible by each $d_{i}$. The graded $k[T_{0},\ldots,T_{N}]/I_{Y}$-module $(k[T_{0},\ldots,T_{N}]/I_{Y})(d)$ is a locally free, coherent rank 1 sheaf of $\mathcal{O}_{Y}$-modules on $Y$. We denote this sheaf by $\mathcal{O}_{Y}(d)$. For each $i$ such that $T_{i}$ is not in the radical of $I_{Y}$, the restriction $(\mathcal{O}_{Y})_{|V_{i}}$ of $\mathcal{O}_{Y}$ to  $V_{i}\ \equiv(\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}])-Z(T_{i}))\cap Y$ is generated by the image of $T_{i}^{d/d_{i}}$ viewed as a global section of $\mathcal{O}_{Y}(d)$.
@@
It suffices to assume that $I_{Y}$ is a radical ideal . Moreover, by considering each irreducible component of $Y$, we may assume that $Y$ is irreducible. Then [8, Ex 9.5] applies to verify that the variey $Z(I_{Y})$ is isomorphic to its image in $\operatorname{Proj}\nolimits S$, where $S\ =k[T_{0},\ldots,T_{N}]_{(d)}\subset k[T_{0},\ldots,T_{N}]$ is the graded subalgebra generated by (weighted) homogeneous polynomials each of whose weights is divisible by $d$ and replacing $I_{Y}$ by the homogenous ideal in $S$ defining the image of $Y$ in $\operatorname{Proj}\nolimits S$. Thus, replacing $k[T_{0},\ldots,T_{N}]$ by $S$ with grading modified by dividing degrees by $d$, we may apply [16, Prop II.5.12] (which is stated for graded submodules of a graded commutative algebra generated in degree 1). ∎
@@
We next present the projectivization of the universal operator $\Theta_{\mathbb{G},r}$. This operator has been used in [14] to produce global invariants of $\mathbb{G}$-modules (especially, non-trivial vector bundles on $\mathbb{P}\mathbb{G}_{(r)}$) in contrast to the Jordan type function $JT_{\mathbb{G},r,M}(-)$ which is a “pointwise on $\mathbb{P}\mathbb{G}_{(r)}$” invariant of $M$.
@@
The universal operator $\Theta_{\mathbb{G},r}$ of Theorem 2.11 determines a global section
$\tilde{\Theta}_{\mathbb{G},r}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
@@
    For any non-zero divisior $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$, the natural isomorphism of $(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})})_{U_{h}}$-modules
@@
    If $\overline{x}$ is a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the condition that $h(\overline{x})\not=0$ and if $x$ is a scheme theoretic point of $V_{r}(\mathbb{G})$ projecting to $\overline{x}$, then the restriction $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathbb{G},r}$ is given by
@@
The first assertion of the theorem follows from Theorem 2.11(3) and the correspondence between graded $k[V_{r}(\mathbb{G})]$-modules and quasi-coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$. The second assertion follows from the standard description of principal open subsets of $\operatorname{Proj}\nolimits R$ for a graded ring $R$ in terms of the degree 0 part of $R[h^{-1}]$ for homogenous elements $h\in R$.

### v2

@@ 2. Jordan type functions for $\mathbb{G}$-modules
(2.12.2) $(\mathcal{U}_{\mathbb{G},r})_{*}:k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\quad\to\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$
the map of group algebras over $k[V_{r}(\mathbb{G})]$ induced by $\mathcal{U}_{\mathbb{G},r}$. In other words, $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the $k[V_{r}(\mathbb{G})]$-dual of the map on coordinate algebras $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ given by $\mathcal{U}_{\mathbb{G},r}$.
By [16, Prop 2.9], a closed embedding $i:\mathbb{H}\hookrightarrow\mathbb{G}$ of affine group schemes of finite type over $k$ determines the commutative square
(2.12.3) ${\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{H},r})_{*}}$${\mathbb{H}_{(r),k[V_{r}(\mathbb{H})]}].}$
We recall and reformulate the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}$ for $\mathbb{G}_{(r)}$-modules introduced in [16, Defn 2.1].
@@
where $(\mathcal{U}_{\mathbb{G},r})_{*}$ is defined in (2.12.2).
In [16, Ex 2.6], the reader will find a few explicit computations of $\Theta_{\mathbb{G},r}$. We mention two such examples consistent with the indexing used in this paper.
@@
Consider $\mathbb{G}=\mathbb{G}_{a}^{\times s}$ whose Lie algebra is the dimension $s$ Lie algebra $\mathfrak{g}_{a}^{\oplus s}$ (which is commutative, with $p$-th power map trivial). Identify $V_{1}(\mathbb{G})$ with $\mathbb{A}^{r}=\operatorname{Spec}\nolimits k[t_{0},\ldots,t_{s-1}]$ and identify the group algebra $k\mathbb{G}_{(1)}$ with $k[x_{0},\ldots,x_{s-1}]/(x_{i}^{p})\\ =\ \mathfrak{u}(\mathfrak{g}_{a}^{\oplus s})$, the restricted enveloping algebra of the restricted Lie algebra $\mathfrak{g}_{a}^{\oplus s}$. Then
$\Theta_{\mathbb{G},1}\ =\ \sum_{i=0}^{r-1}t_{i}\otimes x_{i}\ \in\ k[V_{1}(\mathbb{G}_{a}]\otimes\mathfrak{u}(\mathfrak{g}_{a}^{\oplus s}).$
Consider $\mathbb{G}=\mathbb{G}_{a}$ whose Lie algebra is $\mathfrak{g}_{a}$ and some $r>0$. Identify $V_{r}(\mathbb{G}_{a})$ with $\operatorname{Spec}\nolimits k[T_{0},\ldots,T_{r-1}]$ with $T^{i}$ given grading $p^{i}$ and identify the group algebra $k\mathbb{G}_{(r)}$ with $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. Then
$\Theta_{\mathbb{G}.r}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ +\ \cdots\quad\in\ k[V_{r}(\mathbb{G}_{a})]\otimes k\mathbb{G}_{a(r)},$
where the non-explicit, remaining terms of $\Theta_{\mathbb{G}_{a},r}$ each consist of a polynomial homogeneous of weighted degree $p^{r-1}$ in the $T_{i}$’s tensor a monomial in the $u_{i}$’s of degree at least 2. (See [27, Eqn 6.5.1].)
@@
    Let $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ be a closed subgroup scheme determining $\phi^{*}:k[V_{r}(\mathbb{G})]\twoheadrightarrow k[V_{r}(\mathbb{H})]$ and $\phi_{(r)*}:k\mathbb{H}_{(r)}\hookrightarrow k\mathbb{G}_{(r)}$. Then the restriction of $\Theta_{\mathbb{G},r}$ along $\phi^{*}$, $(\phi^{*}\otimes 1)(\Theta_{\mathbb{G},r})\in k[V_{r}(\mathbb{H})]\otimes k\mathbb{G}_{(r)}$, equals the image of $(1\otimes\phi_{(r)*})(\Theta_{\mathbb{H},r})$.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, $\Theta_{\mathbb{G},r}$ determines a $p$-nilpotent, $k[V_{r}(\mathbb{G})]$-linear operator $\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\ \to\ k[V_{r}(\mathbb{G})]\otimes M$.
@@
Consequently, the restriction of $\Theta_{\mathbb{G},r}=(\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})$.
The naturality of $\mathbb{G}\mapsto\Theta_{\mathbb{G},r}$ with respect to closed embeddings $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ as in assertion (2) follows from the commutativity of (2.12.3).
@@
The fact that this action is $p$-nilpotent follows from the observation that the $i$-th power of $\Theta_{\mathbb{G},r}$ equals $(\mathcal{U}_{\mathbb{G},r})_{*}((1\otimes u_{r-1})^{i})$. Assertion (4) follows directly from assertion (1).
The last assertion is stated and proved in [16, Prop 2.11]. The proof proceeds by first observing that it suffices to verify that the restriction of the $k[V_{r}(\mathbb{G})]$-linear map $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ to $1\otimes k[\mathbb{G}_{(r)}]$ reads off the coefficient of $t^{p^{r-1}}$ in $k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\simeq k[V_{r}(\mathbb{G})]\otimes k[t]/t^{p^{r}}$. The authors then verify that the actions of $\mathbb{A}^{1}$ on $\mathbb{G}_{a(r)}$ and $V_{r}(\mathbb{G})$ are appropriately compatible. ∎
Definition 2.16.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, we define the function
(2.16.1) $JT_{\mathbb{G},r,M}(-):V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}$
from the set of scheme-theoretic points of $V_{r}(\mathbb{G})$ to the set of points of the partially orders set $\mathcal{Y}$ by sending $x\in V_{r}(\mathbb{G})$ to
$JT_{\mathbb{G},r,M}(x)\quad\equiv\quad JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)),$
the Jordan type of the $k(x)[u]/u^{p}$-module obtained by restricting $k[V_{r}(\mathbb{G})]\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$; equivalently, $JT_{\mathbb{G},r,M}(x)$ is the Jordan type of $(\mu_{x})_{*}(1\otimes u_{r-1})$ acting on $k(x)\otimes M$.
By Theorem 2.15(1), $JT_{\mathbb{G},r,M}(x)$ equals the Jordan type of $k(x)\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r,M}\ \in\ End_{k(x)}(k(x)\otimes M)$.
Let $R$ be a graded, commutative $k$-algebra; for example, $R\ =\ k[V_{r}(\mathbb{G})]$. For any graded $R$-module $M_{*}$, we denote by $(M_{*})^{\sim}$ the associated quasi-coherent sheaf on $\operatorname{Proj}\nolimits(R)$. We shall often consider graded $R$-modules of the form $R\otimes V$ where $V$ is a $k$-vector space and $R\otimes V$ inherits the grading of $R$.
If $M_{*}$ is a graded $R$-module, we denote by $M_{*}(n)$ the graded module obtained from $M_{*}$ by “shifting down by nn” the degrees of $M_{*}$; thus, $(M_{*}(n))_{d}=M_{n+d}$. We use $(M_{*})^{\sim}(n)$ to denote $(M_{*}(n))^{\sim}$. The operator $\Theta_{\mathbb{G},r,M}$ of Theorem 2.15(3) multiplies degrees by $p^{r-1}$, so that viewed as an operator on graded modules one should write $\Theta_{\mathbb{G},r,M}$ as
@@
Because $k$ has positive characteristic, the informative survey by M. Beltrametti and L. Robbiano in [2] does not apply to our consideration of $\mathbb{P}V_{r}(\mathbb{G})$. Because $V_{r}(\mathbb{G})$ is usually not generated by homogeneous elements of degree 1, we supplement the clear exposition of R. Hartshorne in [19, §II.5] with [10, Ex 9.5] when considering the relationship between graded modules over a commutative, graded $k$-algebra $R$ and sheaves of modules on $\operatorname{Proj}\nolimits(R)$.
@@
Consider a polynomial algebra $k[T_{0},\ldots,T_{N}]$ with each $T_{i}$ given a “weighted” degree $d_{i}$. Assume that $I_{Y}\subset k[T_{0},\ldots,T_{N}]$ is an ideal generated by polynomials $F_{j}(\underline{T})$ which are (weighted) homogeneous. Then, the zero locus $Y=Z(I_{Y})\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$ is a weighted projective variety associated to the commutative, graded algebra $k[T_{0},\ldots,T_{N}]/I_{Y}$.
Let $d$ be a positive integer divisible by each $d_{i}$. The graded $k[T_{0},\ldots,T_{N}]/I_{Y}$-module $(k[T_{0},\ldots,T_{N}]/I_{Y})(d)$ is a locally free, coherent rank 1 sheaf of $\mathcal{O}_{Y}$-modules on $Y$. We denote this sheaf by $\mathcal{O}_{Y}(d)$. For each $i$ such that $T_{i}$ is not in the radical of $I_{Y}$, the restriction $(\mathcal{O}_{Y}(d))_{|V_{i}}$ of $\mathcal{O}_{Y}(d)$ to open subvariety
$V_{i}\ \equiv(\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}])-Z(T_{i}))\cap Y\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$
is generated as a free, rank 1 $\mathcal{O}_{V_{i}}$-module by the image of $T_{i}^{d/d_{i}}\in\mathcal{O}_{Y}(d)(Y)$ in $\mathcal{O}_{Y}(d)(V_{i})$.
@@
It suffices to assume that $I_{Y}$ is a radical ideal . Moreover, by considering each irreducible component of $Y$, we may assume that $Y$ is irreducible. Then [10, Ex 9.5] applies to verify that the variety $Z(I_{Y})$ is isomorphic to its image in $\operatorname{Proj}\nolimits S$, where $S\ =k[T_{0},\ldots,T_{N}]_{(d)}\subset k[T_{0},\ldots,T_{N}]$ is the graded subalgebra generated by (weighted) homogeneous polynomials each of whose weights is divisible by $d$ and the image of $Z(Y)$ is given by the homogeneous ideal $I_{Y}\cap S\subset S$ which is generated by elements of degree $d$. Thus, replacing $k[T_{0},\ldots,T_{N}]$ by $S$ with grading modified by dividing degrees by $d$, we may apply [19, Prop II.5.12] (which is stated for graded submodules of a commutative, graded algebra generated in degree 1). ∎
@@
We next present the projectivization of the universal operator $\Theta_{\mathbb{G},r}$. This operator has been used in [16] to produce global invariants on $\mathbb{P}V_{r}(\mathbb{G})$ for $\mathbb{G}$-modules (especially, non-trivial vector bundles on $\mathbb{P}V_{r}(\mathbb{G})$) in contrast to the Jordan type function $JT_{\mathbb{G},r,M}(-)$ which is a “point-wise invariant” on $\mathbb{P}V_{r}(\mathbb{G})$ invariant for a $\mathbb{G}$-module $M$. See Section 6 below.
@@
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. The universal operator $\Theta_{\mathbb{G},r}$ of Theorem 2.15 determines a global section
$\tilde{\Theta}_{\mathbb{G},r}\quad\in\quad\Gamma(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
@@
    For any non-zero divisior $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[V_{r}(\mathbb{G})][h^{-1}])_{0}\ \subset\mathbb{P}V_{r}(\mathbb{G})$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})})_{U_{h}}$-modules
@@
    If $\overline{x}$ is a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the condition that $h(\overline{x})\not=0$ and if $x$ is a scheme theoretic point of $V_{r}(\mathbb{G}){\text{-}}\{0\}$ projecting to $\overline{x}$, then the restriction $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathbb{G},r}$ is given by
@@
The first assertion of the theorem follows from Theorem 2.15(3) and the correspondence between graded $k[V_{r}(\mathbb{G})]$-modules and quasi-coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$. The second assertion follows from the standard description of principal open subsets of $\operatorname{Proj}\nolimits R$ associated to homogenous elements $h\in R$ as $\operatorname{Spec}\nolimits R[h^{-1}]$.